OpenStudy (anonymous):

Find the Projection of v onto u where u=(2/3)i-(2/3)j-(1/3)k and v=2i-2j+2k

OpenStudy (anonymous):

The projection of a vector \(\vec{a}\) onto another vector \(\vec{b}\) is given by \[\|a\|\cos\theta\times\frac{\vec{b}}{\|\vec{b}\|}\] where \(\theta\) is the angle between the vectors. Recall that \[\vec{a}\cdot\vec{b}=\|a\|\|b\|\cos\theta~~\iff~~\cos\theta=\frac{\vec{a}\cdot\vec{b}}{\|a\|\|b\|}\]

OpenStudy (anonymous):

All this means you need to find the dot product of the given vectors along with the norms of each vector.

OpenStudy (anonymous):

Would the opposite also be true?

OpenStudy (anonymous):

projection of B onto A

OpenStudy (anonymous):

By "opposite" you must mean switch \(a\) with \(b\) in the formulas above? Then yes. The projection of \(b\) onto \(a\) would be \(\|\vec{b}\|\cos\theta\dfrac{\vec{a}}{\|\vec{a}\|}\).

OpenStudy (anonymous):

\[\theta=\cos^{-1} \frac{ UV \ }{ \left| UV \right|^2 }V\]

OpenStudy (anonymous):

oh alright thank you

OpenStudy (anonymous):

yw